Suppose T in End(F^2) is defined by
T(w,z)=(-z,w)
Find the eigenvalues and eigenvectors of T if F=R.
1
Expert's answer
2020-06-11T16:45:53-0400

The standard basis for R2 is (1,0) and (0,1)

T(1,0)=(0,1)

T(0,1)= (-1,0)

Thematrix of the given linear transformation is given by

A= "\begin{bmatrix}
 0 & -1 \\
 1 & 0
\end{bmatrix}"

The characteristic polynomial is given by

"Det(A-xI) = \begin{vmatrix}
 -x& -1\\
 1& -x
\end{vmatrix}" =0


"x^2 +1 =0\\"

But if F=R, then the characteristic polynomial has no zeroes in R.

So, there are no eigenvalues and no eigenvectors when F=R.